Standard Direct Materials Cost per Unit from Variance Data

The following data relating to direct materials cost for October of the current year are taken from the records of Good Clean Fun Inc., a manufacturer of organic toys:

Quantity of direct materials used

7,300 lbs.

Actual unit price of direct materials

$2.2 per lb.

Units of finished product manufactured

1,755 units

Standard direct materials per unit of finished product

4 lbs.

Direct materials quantity variance—unfavorable

$672

Direct materials price variance—favorable

$1,460

Determine the standard direct materials cost per unit of finished product, assuming that there was no inventory of work in process at either the beginning or the end of the month. If required, round your standard cost per unit answer to two decimal places.
Answer #1

Total Direct material variance = (1460-672) = 788 Favorable

Total material variance = Standard cost-actual Cost

788 = (4*X*1755)-(7300*2.2)

788 = 7020X -16060

-7020X = -16848

X = 2.40

Standard direct material cost per unit of finished product = 4*2.4 = 9.60